A cantilever beam of cross-section 90 mm. width 120 mm deep carries a UDL of 12 KN/m. over the entire length and a concentrated load of 15 KN at the right end. Find the bending stress in the beam, when the length of beam is 10 m.


1
Expert's answer
2021-12-21T02:11:03-0500

Length of beam = 2000 mm

Load at failure = 2 kN

Section dimensions = 40 mm x 60 mm

Moment if inertia

I = bd3/12 = (40) (603) / 12 = 7.2 x 105 mm4

Bending moment about fixed end

M = WL = (2) (2) = 4 kN/m

Bending stress

M / l = "\sigma" / y

y = depth / 2 = 60 / 2 = 30 mm

"\sigma" = 166.67 N/mm2